OpenStudy (anonymous):

whats the product of (x+3)(x-3)?

OpenStudy (bahrom7893):

this is a difference of the squares

OpenStudy (bahrom7893):

\[(a-b)(a+b)=a^2-b^2\]

OpenStudy (bahrom7893):

a is x, b is 3. Plug them into the formula

sam (.sam.):

(x+3)(x-3) Multiply each term in the first group by each term in the second group using the FOIL method. FOIL stands for First Outer Inner Last, and is a method of multiplying two binomials. First, multiply the first two terms in each binomial group. Next, multiply the outer terms in each group, followed by the inner terms. Finally, multiply the last two terms in each group. (x*x+x*-3+3*x+3*-3) Simplify the FOIL expression by multiplying and combining all like terms. (x^(2)-9)

OpenStudy (bahrom7893):

once u memorize the square formulas this will become a joke.

OpenStudy (anonymous):

\[(a−b)(a+b)=a ^{2}−b ^{2}\] is the formula \[x^{2}-3^{2}\]